# sample.bib

### v4

@@ A.2 Standard benchmarks
Appendix B Action Associative Retrieval
@@
There are two states in this environment: $S_{0}$ and $S_{1}$. The agent appears in state $S_{0}$ and by performing the action $a_{0}\in\{0,1\}$ moves to state $S_{1}$. Next, the agent must take $N-2$ steps to move from state $S_{1}$ to state $S_{1}$ by performing action $a=2$ (no op.). At the end of the episode, the agent must perform the same action that moved it from state $S_{0}$ to state $S_{1}$ in order to move from state $S_{1}$ to state $S_{0}$. Thus, the action $a\in\{0,1,2\}$. Agent observations $o=[state,flag,noise]$, where $state\in\{0,1\}$ is the index of the current state, $flag\in\{0,1\}$ is a flag equal to $1$ in case the next step requires returning to the initial state and equal to 00 otherwise, $noise\in\{-1,0,+1\}$ is the noise channel. The agent receives a $+1$ reward if it returns to the initial state $S_{0}$ by performing the action that took it out from the $S_{0}$ to the $S_{1}$, and $-1$ in other cases.
Appendix C Training
@@
To train RATE on this environment, we collected dataset of 5000 trajectories of length 90 environment steps using a pre-trained A2C agent (trained without disappearing oh the pillar) [DBLP:journals/corr/abs-1904-01806]. Since the pillar disappears at time $t=45$, all trajectories start at time $t=0$ and end at time $t=90$ so that the information about the color of the pillar is guaranteed to be used in training. In this experiment, we compared DT with context length $K=90$ to our RATE model with context length $K=30$ and partitioning the trajectory into $N=3$ segments. Thus, when trained, RATE also handles sequences of length $90$, since its effective context is $K_{eff}=N\times K=90$, but only processed subsequences of length $K=30$.
@@
The model names are written in the format MODEL-N, where $N$ is the number of segments of length $K=30$ into which the training trajectories can be partitioned. Thus, DT-3 was trained on trajectories of length $T\leq 3\times 30=90$, and DT-5 was trained on trajectories of length $T\leq 5\times 30=150$ with context lengths $K=90$ and $K=150$, respectively. RATE-3 and RATE-5 were trained on similar trajectories as DT-3 and DT-5, but with each trajectory divided into $3$ and $5$ segments, respectively, during training, enabling the training of a model with a context length of $K=30$ on trajectories of length $T=90$ and $T=150$. On this environment, RATE was trained using curriculum learning because this approach produces better results. However, it is important to note that the T-Maze task is successfully solved by the RATE model without using curriculum learning (see section 5 for details). All the trajectories used in training start from $t=0$, i.e., from the moment of receiving a clue. The work showed that the use of Curriculum Learning (CL) allows one to achieve significantly better results when solving the T-Maze environment. In this approach, the model is first trained on the set of trajectories $Q_{1}=q_{1}$ of length $K\times 1$, then the trained model is re-trained on the set of trajectories $Q_{2}=q_{1}\cup q_{2}$, where the set $q_{2}$ is formed by trajectories of length $K\times 2$, and so on (in order of increasing complexity of the trajectories). Thus, for the $N$ segments considered during training, the set $Q_{N}=\bigcup_{i=1}^{N}q_{i}$ is used.
@@
To train DT and RATE on Memory Maze, the same approach was used as for ViZDoom-Two-Colors environment, except that trajectories were sampled not from $t=0$ but from $t:\sum_{t^{\prime}=t}^{t+90}r_{t^{\prime}}\geq 2$. This is because in the [pasukonis2022memmaze] offline training data posted by the authors, the average value of the total reward over all trajectories where the total reward is $>0$ is $1.06$. Furthermore, the trajectories themselves are not collected by an agent with an optimal or suboptimal policy, but by an agent navigating to a random object in the maze. As in the ViZDoom-Two-Colors case, training for DT was performed with a context length of $K=90$ and for RATE with a context length of $K=30$ and number of segments $N=3$, i.e., effective context length $K_{eff}=N\times K=3\times 30=90$.
Minigrid.Memory.
To train RATE and DT on this environment, trajectories were sampled in the same manner as for T-Maze. An environment configuration with a maze of size $13\times 13$ was used as a basis, and trajectories were collected using PPO [schulman2017proximal] with TransformerXL [pleines2023trxlppo] with a context length equal to the maximum episode duration. Since the maximum episode duration is $95$, training proceeded in the following setting: for DT the context length $K=30$, for RATE the context length $K=10$ and the number of segments $N=3$. All trajectories, as in T-Maze, are sampled from time $t=0$.
@@
When training RATE on Atari games and MuJoCo control tasks, sequences of length $T=90$ (Atari) and $T=60$ (MuJoCo) were sampled randomly from the original trajectories in the dataset. These trajectories were then divided into $N=3$ segments of length $K=30$ (Atari) and $K=20$ (MuJoCo), forming an effective context of length $K_{eff}=N\times K=90$ ($60$ for the MuJoCo).
For the Atari, we used the identical experimental design described in the DT paper [chen2021decision]. It is worth noting that we presented raw scores for Atari, rather than gamer-normalized scores as described in the DT paper. Table 1 shows the results for Atari environments. RATE outperforms DT significantly in environments like Breakout and Qbert. We attribute this to the observation that, although these environments do not explicitly demand memory, intricate dynamics from the past exert a greater influence on agent behavior than in environments such as SeaQuest. Actions executed in the past notably alter the present state of the environment in Breakout and Qbert, whereas in SeaQuest, such actions hold little significance. For instance, the emergence of enemies and divers in SeaQuest is entirely independent of the agent’s prior actions.
For the MuJoCo, our findings suggest that the conventional strategy of utilizing return is not suitable for our segment-based scheme. The issue arises during the trajectory, where the agent’s return persistently diminishes. However, the true value of the agent’s state at the onset and conclusion of the episode could remain unchanged, provided the agent’s policy performs consistently well. To rectify this discrepancy, we propose a novel evaluation strategy for MuJoCo tasks. In this approach, each segment commences with the maximum return, simulating the scenario where the agent initiates the trajectory anew. This method effectively mitigates the aforementioned issue, enhancing the accuracy of our evaluation process. Our MuJoCo experiments in Table 2 show that this benefits performance significantly for some environments. Thus, using RATE allowed us to obtain the best metrics for MuJoCo in 5/9 cases compared to the other baselines. RATE also outperforms DT in 9/9 tasks.
Appendix D Hyperparameters
In this section, Table 5 lists the hyperparameters used to train the RATE model. Inside the Memory Retention Valve, the number of attention heads $=2$ was used for all environments (with the exception of Memory Maze, for which this parameter is set to $4$).
@@ Appendix D Hyperparameters
Hyperparameter
ViZDoom2C /
Memory Maze
T-Maze /
Minigrid.Memory
Atari MuJoCo
Number of layers 6 8 6 3
Number of attention heads 8 10 8 1
Embedding dimension 128 64 128 128
Context length K 30 30 / 10 30 20
Number of segments 3 3 3 3
Hidden dropout 0.2 0.05 / 0.2 0.2 0.2
Attention dropout 0.05 0 / 0.05 0.05 0.05
Number of memory tokens 15 5 15 15
Max epochs 1000 250 10 10
Batch size 128 64 128 4096
Weight decay 0.1 1e-4 0.1 0.1
Loss function CE CE CE MSE
Optimizer AdamW AdamW AdamW AdamW
MRV activation ReLU ReLU ReLU ReLU
MRV number of attention heads 2 / 4 2 2 2
@@
Minigrid.Memory Linear Conv2D $\times$ 3 (32, 64, 64) / (8, 4, 3) / 0 Embedding
@@
Appendix E Memory Retention Valve
The MRV-G scheme was inspired by the gating mechanism from GTrXL [parisotto2020stabilizing] and implemented as follows:
@@ Appendix E Memory Retention Valve
$z=\sigma(M_{n}W_{z}+M_{n+1}U_{z}-bias)$ (9)
$h=tanh(M_{n}W_{g}+(M_{n+1}\times r)U_{r})$ (10)
@@
The results of the RATE (trained on corridor lengths of $\leq 150$) inference on the T-Maze environment with these MRV configurations are shown in Figure 8 and in Table 3.
@@
Figure 9: Memory Retention Valve configurations used in the ablation study. MRV-CA-2: cross-attention-based MRV which uses an attention mechanism to control the updating of memory embeddings and which is used in the work as the main mechanism. MRV-CA-1: uses the same mechanism as MRV-CA-2 but the updated memory embeddings $M_{n+1}$ are fed to Query, and the incoming memory embeddings $M_{n}$ are fed to Key and Value. MRV-G: gated MRV which uses a gating mechanism similar to the one used in Gated Transformer-XL [parisotto2020stabilizing]. MRV-GRU: uses a GRU [gru] block to process updated memory embeddings with hidden states. MRV-LSTM: uses a LSTM [hochreiter1997long] block to process updated memory embeddings with cached states.
Appendix F Technical details
The Table 7 shows the technical parameters of the training models. Note that the difference between the number of DT and RATE parameters is small and equal to $\delta p$ = d_model $\times$ num_mem_tokens $\sim 10^{3}$. Training RATE with trajectory splitting into $N$ segments allows $\sim N$ smaller GPU memory size usage than for DT. The training was conducted using a single NVIDIA Tesla A100 80 Gb graphics card.
The code is available at https://github.com/AIRI-Institute/RATE.

### v5

@@ B.1 Memory-intensive environments
Memory Maze.
In this first-person view 3D environment (Pasukonis et al., 2022), the agent appears in a randomly generated maze containing several objects of different colors at random locations. The agent’s task is to find an object of the same color in the maze as the outline around its observation image. After the agent finds an object of the desired color and steps on it, the color of the outline changes and the agent must find another object. The agent receives a $+1$ reward for stepping on the correct object. Otherwise, it receives no reward. The duration of an episode is a fixed number and is equal to $1000$. Thus, the agent’s task is to find as many objects of the desired color as possible in a limited time. The agent’s effectiveness in this environment depends on its ability to memorize the structure of the maze and the location of objects in it in order to find the desired objects faster. Using the Dreamer model (Hafner et al., 2019) to collect dataset of $5000$ trajectories only achieved an average award of $4.7$ per episode, i.e., a rather sparse dataset.
Minigrid-Memory.
Minigrid-Memory (Chevalier-Boisvert et al., 2023) is a 2D grid environment designed to test an agent’s long-term memory and credit-assignment (Ni et al., 2023). The environment map is a T-shaped maze with a small room with an object inside it at the beginning of the corridor. The agent appears at a random coordinate in the corridor. The agent’s task is to reach the room with the object and memorize it, then reach the junction at the end of the maze and make a turn in the direction where the same object is located as in the room at the beginning of the maze. A reward $r=1-0.9\times\frac{t}{T}$ is given for success, and 00 for failure. The episode ends after any agent turns at a junction or after a limited amount of time (95 steps) has elapsed. The agent’s observations are limited to a $3\times 3$ size frame. $10000$ trajectories with grid size $31$x31313131 were collected using PPO (Schulman et al., 2017) with TransformerXL (Pleines et al., 2023) with a context length equal to the maximum episode duration.
B.2 Standard benchmarks
Atari games.
For the Atari game environments (Bellemare et al., 2013), we used the same dataset as in DT, namely the DQN replay dataset with grayscale state images (Agarwal et al., 2020). This dataset contains $500$ thousand of the $50$ million steps of an online DQN (Mnih, 2013) agent for each game. We use the following set of games: SeaQuest, Breakout, Pong and Qbert.
MuJoCo.
Despite the fact that memory is not required in decision making in control environments like MuJoCo (Fu et al., 2021), we conducted additional experiments in this environment to compare with DT. For the continuous control tasks, we selected a standard MuJoCo locomotion environment and a set of trajectories from the D4RL benchmark (Fu et al., 2021). Since we chose DT and TAP as the main models for comparison on this data, we focused on the environments used in both works (HalfCheetah, Hopper, and Walker). We used three different dataset settings: 1) Medium1111 million timesteps generated by a “medium” policy that achieves about a third of the score of an expert policy; 2) Medium-Replay – the replay buffer of an agent trained with the performance of a medium policy (about $200$k–$400$k timesteps in our environments); 3) Medium-Expert1111 million timesteps generated by the medium policy concatenated with 1 million timesteps generated by an expert policy. The scores for the MuJoCo experiments are normalized such that 100 represents an expert policy, following the benchmark protocol outlined in (Fu et al., 2021). The performance metrics for Conservative Q-Learning (CQL) and Trajectory Autoencoding Planner (TAP) are reported from the TAP paper (Jiang et al., 2023), and for DT from the DT paper (Chen et al., 2021), as they use the same dataset and evaluation protocol.
Figure 8: Action Associative Retrieval.
Appendix C Action Associative Retrieval
As shown in subsection 4.2, DT has a SR $=50\%$ for inference at corridor lengths longer than the transformer context length. This is due to the fact that even a DT trained on balanced data has a slight bias in the predicted probability towards one of the two required actions, which leads to the fact that when $t>K$ the agent constantly produces only one action: up or down. In turn, the presence of memory in the agent allows us to combat this problem.
To check how the agent’s performance changes during training, we design an Action Associative Retrieval (AARFigure 8 environment.
@@ Appendix C Action Associative Retrieval
There are two states in this environment: $S_{0}$ and $S_{1}$. The agent appears in state $S_{0}$ and by performing the action $a_{0}\in\{0,1\}$ moves to state $S_{1}$. Next, the agent must take $N-2$ steps to move from state $S_{1}$ to state $S_{1}$ by performing action $a=2$ (no op.). At the end of the episode, the agent must perform the same action that moved it from state $S_{0}$ to state $S_{1}$ in order to move from state $S_{1}$ to state $S_{0}$. Thus, the action $a\in\{0,1,2\}$. Agent observations $o=[state,flag,noise]$, where $state\in\{0,1\}$ is the index of the current state, $flag\in\{0,1\}$ is a flag equal to $1$ in case the next step requires returning to the initial state and equal to 00 otherwise, $noise\in\{-1,0,+1\}$ is the noise channel. The agent receives a $+1$ reward if it returns to the initial state $S_{0}$ by performing the action that took it out from the $S_{0}$ to the $S_{1}$, and $-1$ in other cases. The training dataset consists of oracle-generated $6000$ trajectories with positive reward.
This condition means that if the agent has memory, the sum of the average conditional probabilities over all experiments will be greater than one, i.e., these probabilities are independent of each other. Provided that the sum of these probabilities is less than or equal to one, the agent will choose at best the same target action in most experiments, even if another action is required.
where $a^{0},a^{1}\in\mathcal{A}$ – two mutually exclusive actions leading to a reward; $t$ is the step at which the final action is required; $N_{0},N_{1}$ are the number of experiments in environments where target action $a_{t}=a^{0}$ and $a_{t}=a^{1}$, respectively.
In the results Figure 9, the first $1\%$ of training steps was removed because it corresponds to the beginning of the training and is unrepresentative. Blue dots correspond to the beginning of training, red dots to the end of training. As can be seen from Figure 9, during training, the probabilities $p_{i}(a_{t}=a^{0}|a_{0}=a^{0})$ and $p_{i}(a_{t}=a^{1}|a_{0}=a^{1})$ on the training trajectories have a strong positive correlation ($R_{train}^{DT}=1.00$ and $R_{train}^{RATE}=0.97$), where $R$ – correlation coefficient. This indicates that within-context (effective context) DT and RATE models are able to predict both $a^{0}$ and $a^{1}$ actions equally well.
At the same time, during validation, for the RATE model this pattern is preserved – the red points corresponding to the probabilities of choosing actions $a^{0}$ and $a^{1}$ are in the upper right part of the graph, positive correlation persists ($R_{val}^{RATE}=0.80$). On the other hand, in the DT case, the cluster of red dots is skewed toward choosing action $a^{1}$ and action $a^{0}$ with equal probabilities equal to $0.5$. Thus, in sum, these probabilities are less or equal to one, as evidenced by a strong negative correlation ($R_{val}^{DT}=-0.97$). The results confirm the inability of DT to generalize on trajectories whose lengths exceed the context length and the ability of RATE to handle such tasks.
Appendix D Training
This section provides additional details on the training process of the baselines considered in the paper. It is important to note that when training RATE in the transformer decoder the feed-forward network block was disabled, because without it on some environments the training results are slightly better. However, other transformer-based baselines were trained with the standard transformer decoder.
@@ Appendix D Training
Since the pillar disappears at time $t=45$, all trajectories start at time $t=0$ and end at time $t=90$ so that the information about the color of the pillar is guaranteed to be used in training. In this experiment, we compared DT with context length $K=90$ to RATE, RMT, and TrXL models with context length $K=30$ and partitioning the trajectory into $N=3$ segments. Thus, when trained, RATE also handles sequences of length $90$, since its effective context is $K_{eff}=N\times K=90$, but only processed subsequences of length $K=30$.
@@
The model names are written in the format MODEL-N, where $N$ is the number of segments of length $K=30$ into which the training trajectories can be partitioned. Thus, DT-3 was trained on trajectories of length $T\leq 3\times 30=90$ with context length. RATE-3 was trained on similar trajectories as DT-3, but with each trajectory divided into $3$ segments, during training, enabling the training of a model with a context length of $K=30$ on trajectories of length $T=90$. All the trajectories used in training start from $t=0$, i.e., from the moment of receiving a clue.
@@
To train RATE, DT, RMT, and TrXL on Memory Maze, the same approach was used as for ViZDoom-Two-Colors environment, except that trajectories were sampled not from $t=0$ but from $t:\sum_{t^{\prime}=t}^{t+90}r_{t^{\prime}}\geq 2$.
As in the ViZDoom-Two-Colors case, training for DT was performed with a context length of $K=90$ and for RATE, RMT, and TrXL with a context length of $K=30$ and number of segments $N=3$, i.e., effective context length $K_{eff}=N\times K=3\times 30=90$.
Minigrid-Memory.
To train RATE, DT, RMT, and TrXL in this environment, trajectories were sampled in the same manner as for T-Maze. An environment configuration with a maze of size $31$x31313131 was used as a training configuration. Since the maximum episode duration is $95$, training proceeded in the following setting: for DT the context length $K=30$, for RATE, RMT, and TrXL the context length $K=10$ and the number of segments $N=3$. All trajectories, as in T-Maze, are sampled from time $t=0$.
@@
When training RATE on Atari games and MuJoCo control tasks, sequences of length $T=90$ (Atari) and $T=60$ (MuJoCo) were sampled randomly from the original trajectories in the dataset. These trajectories were then divided into $N=3$ segments of length $K=30$ (Atari) and $K=20$ (MuJoCo), forming an effective context of length $K_{eff}=N\times K=90$ ($60$ for MuJoCo).
For Atari, we used the identical experimental design described in the DT paper (Chen et al., 2021). It is worth noting that we presented raw scores for Atari, rather than gamer-normalized scores as described in the DT paper. Table 3 shows the results for Atari environments. RATE outperforms DT significantly in environments like Breakout and Qbert. We attribute this to the observation that, although these environments do not explicitly demand memory, intricate dynamics from the past exert a greater influence on agent behavior than in environments such as SeaQuest. Actions executed in the past notably alter the present state of the environment in Breakout and Qbert, whereas in SeaQuest, such actions hold little significance. For instance, the emergence of enemies and divers in SeaQuest is entirely independent of the agent’s prior actions.
For MuJoCo, our findings suggest that the conventional strategy of utilizing return is not suitable for our segment-based scheme. The issue arises during the trajectory, where the agent’s return persistently diminishes. However, the true value of the agent’s state at the onset and conclusion of the episode could remain unchanged, provided the agent’s policy performs consistently well. To rectify this discrepancy, we propose a novel evaluation strategy for MuJoCo tasks. In this approach, each segment commences with the maximum return, simulating the scenario where the agent initiates the trajectory anew. This method effectively mitigates the aforementioned issue, enhancing the accuracy of our evaluation process. Our MuJoCo experiments in Table 4 show that this benefits performance significantly for some environments. Thus, using RATE allowed us to obtain the best metrics for MuJoCo in 4/9 cases compared to the other baselines. RATE also outperforms DT in 9/9 tasks.
@@
Hyperparameter ViZDoom2C / Memory Maze T-Maze / Minigrid-Memory Atari MuJoCo Number of layers 6 8 6 3 Number of attention heads 8 10 8 1 Embedding dimension 128 / 64 64 128 128 Context length K 30 30 / 10 30 20 Number of segments 3 3 3 3 Hidden dropout 0.2 / 0.5 0.05 / 0.2 0.2 0.2 Attention dropout 0.05 / 0.2 0 / 0.05 0.05 0.05 Number of memory tokens 5 / 15 5 / 15 15 15 Number of cached tokens (mem_len) 300 / 360 0 / 180 360 2 Max epochs 100 / 80 50 / 250 10 10 Batch size 64 64 128 4096 Weight decay 0.1 0.1 0.1 0.1 Loss function CE CE CE MSE Optimizer AdamW AdamW AdamW AdamW MRV activation ReLU ReLU ReLU ReLU MRV number of attention heads 2 / 4 4 / 1 2 2 Learning rate 3e-4 3e-4 3e-4 1e-4 AdamW $(\beta_{1},\beta_{2})$ (0.9, 0.95) (0.9, 0.95) (0.9, 0.95) (0.9, 0.95)
Appendix E Results presentation
This section provides information on how the presented experimental results were obtained. $N_{runs}$ denotes the number of model runs; $N_{seeds}$ denotes the number of inference episodes with different seeds; sem denotes standard error of the mean, and std denotes standard deviation.
Table 9: Experimental parameters used to present the final results.
Environment $N_{runs}$ $N_{seeds}$ Metric Notation
ViZDoom-Two-Colors 6 100 Total reward mean $\pm$ sem
T-Maze 10 100 Success Rate mean $\pm$ sem
Memory Maze 3 100 Total reward mean $\pm$ sem
Minigrid-Memory 3 100 Total reward mean $\pm$ sem
Action Associative Retrieval 10 Success Rate mean $\pm$ sem
Atari 3 100 Total reward mean $\pm$ std
MuJoCo 3 100 Total reward mean $\pm$ std
Appendix F Additional ablation studies
[…]
Figure 10: (left) Investigating the RATE memory tokens noise effect in the ViZDoom-Two-Colors. (right) Results of RATE-3 (trained on corridor lengths $\leq 90$) ablation studies in the T-Maze environment. n_head_ca – number of MRV attention heads, num_mem_tokens – number of memory tokens.
Figure 11: Results of RATE ablation studies in the ViZDoom-Two-Colors environment.
F.1 Additional ViZDoom-Two-Colors ablation
[…]
@@ F.1 Additional ViZDoom-Two-Colors ablation
Minigrid-Memory Linear Conv2D $\times$ 3 (32, 64, 64) / (8, 4, 3) / 0 Embedding
@@
Figure 12: Memory Retention Valve configurations used in the ablation study. MRV-CA-2: cross-attention-based MRV which uses an attention mechanism to control the updating of memory embeddings and which is used in the work as the main mechanism. MRV-CA-1: uses the same mechanism as MRV-CA-2 but the updated memory embeddings $M_{n+1}$ are fed to Query, and the incoming memory embeddings $M_{n}$ are fed to Key and Value. MRV-G: gated MRV which uses a gating mechanism similar to the one used in Gated Transformer-XL (Parisotto et al., 2020). MRV-GRU: uses a GRU (Chung et al., 2014) block to process updated memory embeddings with hidden states. MRV-LSTM: uses a LSTM (Hochreiter & Schmidhuber, 1997) block to process updated memory embeddings with cached states.
F.2 Curriculum Learning
Since in the T-Maze environment, the number of actions at the junction relates to the number of actions when moving straight along the corridor as $\frac{1}{L}$ and tends to 00 as $L$ increases, there is a significant imbalance in the agent’s action distribution, which can cause problems when performing rare class (turning actions) prediction. Theoretically, this situation can be remedied through curriculum learning.
Curriculum learning (CL) is a technique in which a model is trained on examples of increasing difficulty. In this approach, the model is first trained on the set of trajectories $Q_{1}=q_{1}$ of length $K\times 1$, then the trained model is re-trained on the set of trajectories $Q_{2}=q_{1}\cup q_{2}$, where the set $q_{2}$ is formed by trajectories of length $K\times 2$, and so on (in order of increasing complexity of the trajectories). Thus, for the $N$ segments considered during training, the set $Q_{N}=\bigcup_{i=1}^{N}q_{i}$ is used.
In the T-Maze environment, DT, RATE, RMT, and TrXL were trained with and without curriculum learning because this approach theoretically produces better results. However, it is important to note that the T-Maze task is successfully solved by the RATE model without using curriculum learning, and even vice versa – its use slightly degraded performance on long corridors. However, with respect to TrXL, the use of CL yielded slightly better results. The work showed that using CL does not achieve significantly better performance on the T-Maze task. The results of using the CL on the T-Maze environment are presented in Figure 13 (left), and the results of applying noise to memory embeddings to assess its importance are presented in Figure 13 (right).
Figure 13: (left). Results with and without the use of curriculum learning and (right) results of replacing RATE memory tokens with white noise at inference in T-Maze.
F.3 Supplemental MRV ablation
One of the options for implementing the memory tokenization gating mechanism was an approach similar to the one proposed in Gated Transforer-XL (GTrXL) (Parisotto et al., 2020) work. Thus, the MRV-G scheme was inspired by the gating mechanism from GTrXL and implemented as follows:
@@ F.3 Supplemental MRV ablation
$z=\sigma(M_{n}W_{z}+M_{n+1}U_{z}-\texttt{bias})$ (3)
$h=\texttt{tanh}(M_{n}W_{g}+(M_{n+1}\times r)U_{r})$ (4)
@@
The results of the RATE (trained on corridor lengths of $\leq 150$) inference on the T-Maze environment with these MRV configurations are shown in Figure 14 and in Table 6. The results presented in Figure 14 confirm the high stability of RATE when using cross-attention-based MRV (MRV-CA-2), as well as the model’s ability to hold important information in memory embeddings when inference on long tasks.
@@
Appendix G Technical details
The Table 11 shows the technical parameters of the training models. Note that the difference between the number of DT and RATE parameters is small and equal to $\delta p$ = d_model $\times$ num_mem_tokens $\sim 10^{3}$. Training RATE with trajectory splitting into $N$ segments allows $\sim N$ smaller GPU memory size usage than for DT. The training was conducted using a single NVIDIA A100 80 Gb graphics card.
The code is available at https://sites.google.com/view/rate-model/.